OpenStudy (anonymous):

Cassy shoots a large marble (Marble A, mass: 0.06 kg) at a smaller marble (Marble B, mass: 0.03 kg) that is sitting still. Marble A was initially moving at a velocity of 0.7 m/s, but after the collision it has a velocity of –0.2 m/s. What is the resulting velocity of marble B after the collision? Be sure to show your work for solving this problem along with the final answer.

OpenStudy (anonymous):

@Loser66 @Destinymasha @Data_LG2 @sammixboo @surjithayer @pooja195 THIS IS THE WORK I SHOWED PLZ CONFIRM

OpenStudy (anonymous):

[p of Marble A + p of Marble B]initial = [p of Marble A + p of Marble B]final [(0.06 kg * 0.7 m/s) + (0.03 kg * 0 m/s)]initial = [(0.06 kg * -0.2 m/s) + (0.03 kg * x m/s)]final 0.42 kg*m/s + 0 kg*m/s = -0.012 kg*m/s + 0.03x kg*m/s 0.432 kg*m/s = 0.03x kg*m/s x = 14.4 m/s

sammixboo (sammixboo):

Sorry I can't help with much physics :(

OpenStudy (anonymous):

lol that's alright :)

OpenStudy (anonymous):

give me a sec.. i'll try to do it on my own

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

.06*.7=.042

OpenStudy (anonymous):

i got a different answer ..close enough though

OpenStudy (anonymous):

oh there you go.. he found the mistake :D

OpenStudy (anonymous):

.042+.012=.03x

OpenStudy (anonymous):

oh ok now i understand...i forgot the extra zero

OpenStudy (anonymous):

Is this how it's supposed to look like: [p of Marble A + p of Marble B]initial = [p of Marble A + p of Marble B]final [(0.06 kg * 0.7 m/s) + (0.03 kg * 0 m/s)]initial = [(0.06 kg * -0.2 m/s) + (0.03 kg * x m/s)]final 0.042 kg*m/s + 0 kg*m/s = -0.012 kg*m/s + 0.03x kg*m/s 0.054 kg*m/s = 0.03x kg*m/s x = 1.8 m/s